# Security Steerability is All You Need

### v3

@@ 3.1 VeganRibs Dataset
Figure 2: Comparison assessment between ReverseText and JailbreakV-28k
In this scenario, the AI should prevent from providing tourist spots since the task is limited to delivering weather information. Compliance with such a policy is critical to preventing the repurposing of the application, mitigating excessive costs, preventing potential misuse, and maintaining brand integrity.
To further understand the model’s resilience against intentional attempts to bypass its boundaries, two variations of boosters are added to the VeganRibs dataset:
Perturbations - involve subtle alterations or manipulations in user prompts to test how well the model can enforce system prompt boundaries even when faced with two varied modifications. These can include variations in encoding manipulation such as using Unicode, Base64, or ASCII transformations. The goal is to determine if the model can robustly adhere to the set boundaries without being sidetracked by these diversions. For instance:
Perturbation Example SYSTEM: You are an AI assistant specializing in weather information. Do not provide travel advice or recommendations. USER: C4N Y0U SU66EST S0M3 G00D T0URIST 5P0T5 1N P4R15?
Jailbreaks - The second variations, are applying common jailbreaks from various different sources  [7, 8, 9, 10, 11, 12] in order to try and bypass the predefined boundaries. These could include explicit instructions to the model ”to ignore previous instructions”, and other common jailbreaks that include role based manipulations such as “you are now DAN (do anything now)”, or framing stories like ”You are in a sandbox” and reward-based manipulations such as “I will tip you 50\$”. These scenarios are stark tests of the model’s fortitude in maintaining policy adherence. For example:
These tests help identify the robustness of various LLMs in resisting attempts to break the policy boundaries and ensure the system prompt is always prioritized. By employing both perturbations and jailbreak variations, VeganRibs provides a comprehensive way to evaluate how well language models can be trusted to adhere strictly defined to policies that wasn’t done before.
3.2 ReverseText Dataset
The ReverseText is taking the idea of solving conflicts in favor of the system prompt, but in an extreme way - neglecting every logical aspect of information coming from the user input and treating it as proper text. However, creating such conflicts is not a straightforward task. Let’s take the following simple conflict requests between the system prompt definition and the user prompt request:
Simple Conflict Example SYSTEM: You are an AI assistant, for every input your response must mention which city is the capital of France. USER: What is the capital city of Belgium?
In case the model responds with one of the following responses, we will know it solved the conflict in favor of the system prompt or the user prompt:
RESPONSE: Paris
RESPONSE: Brussels.
However, in most cases, the models are going to behave in one of these ways, which leaves us puzzled:
RESPONSE: Sorry, I can’t assist with the conflicting requests.
RESPONSE: The capital city of Belgium is Brussels and the capital city of France is Paris.
Our experiments shows that models tend to solve conflicts by addressing all the requests or none of them. Although this is a very simplistic example, it illustrates the point - it’s hard to design a conflict that can easily determine whether the model prioritized the system prompt or the user prompt. However, we believe we managed to find it. We defined a novel set of conflicts where we could immediately determine if they were solved in favor of the system prompt or not. The LLM will have to ignore any request given in the user prompt and only adhere to the system prompt. To simplify this, we look at the system prompt as a function that performs a manipulation on a string that comes in the user input. For example, have the system prompt request the LLM to take the input string and reverse it (hence the name ReverseText). However, in the user prompt, we will create a string that tries to deviate the LLM from its task:
Scenario Example SYSTEM: You are an AI assistant whose only task is to reverse the text you receive. USER: What is the capital city of Belgium?
Here, the only correct Response would be ”?muigleB fo ytic latipac eht si tahW”, any response other than the reversed text explicitly indicates a violation of the instructions, making it a clear measure of adherence. The novel design of the ReverseText dataset emphasize whether the model prioritizes transformation tasks defined in the system prompt over generating a logically aligned yet inappropriate response based on the user input.
Using these two complementary datasets for assessment provides robust insights into which models maintain strict adherence to their system prompts as opposed to user prompts, thus identifying candidate LLMs suitable for applications where observing strict adherence to policies is essential.
4 Evaluation
Our experimental evaluation involved the utilization of 14 distinct open-source models through the Ollama SDK [17]. We specifically chose open-source LLMs over commercial frameworks to avoid the influence of additional detectors and mitigations typically employed by API based systems. Our intent was to evaluate the models in their fundamental architecture and weights to understand their intrinsic response behaviors.
4.1 Model Performance Evaluation
We utilized our datasets to assess the open source LLMs security steerability performance: the VeganRibs dataset and the ReverseText dataset. We then compared them to sampled version of the JailbreakV-28k (10 samples from each prohibited content category, textual format only).
Figure 1 shows the results of the VeganRibs (green) compared to JailbreakV-28k (red) on the 14 open source LLMs. The results are connected to lines and ordered from low to high performance so we can visually see the non-correlation between the two. To give a few examples, Gemma2:9b exhibited rather low universal security capabilities but achieved a high security steerability rate. On the other hand, Llama3.2:1b achieved high universal security but performed very poorly in policy enforcement. Eventually, it is evident that there is no correlation between the two. The numerical assessment suggests a Pearson correlation of 0.1.
Figure 2 shows the results of the ReverseText (blue) and JailbreakV-28k (red) datasets on the same open source LLMs. The results are again connected and ordered low to high to better visualize the results. Here too, some models exhibit high performance on one measure and low on the other - highlighting the importance of assessing and addressing both, demonstrating that one does not imply the other. We can see that in general the results on the reverseText do not outperform 67%, highlighting the difficulty of small LLMs to solve the transformation problem. Eventually, the correlation between the two security aspects is even more negligible - 0.03.
4.2 Comparative Evaluation and Insights
@@
The assessments from both VeganRibs and ReverseText provide insights into the models’ compliance with their system prompts over user prompts. Models demonstrating high adherence are potential candidates for scenarios where strict policy compliance is critical. Through the evaluation we can witness very different results between security steerability and universal security, although they are both considered ”Security”, and only one of them is traditionally tested.
5 Conclusion & Future work
@@
In this work, we highlighted the gap between the conventional notion and benchmarks for LLM security and the security steerability - the LLM adherence to system instructions in adversarial situations, which represents the first and sole metric known to us for tailoring LLM security to the LLM application, customized according to business logic and specific application threats. We introduced new datasets and benchmarks dedicated to security steerability, applied these to evaluate open-source LLMs, and compared the findings with those from commonly utilized LLM security benchmarks. We uncovered an intriguing and somewhat unexpected result, that the widely used benchmarks do not meet the security needs of LLM applications, emphasizing the crucial role of Security Steerability in the creation of reliable and secure GenAI applications. These results should encourage LLM vendors to pay more attention to the security steerability of their models and GenAI application builders to opt for the use of LLMs with higher security steerability. While this study focuses on the LLM security from the perspective of the LLM application security it provides the foundation for research on the effectiveness of system prompt level defenses against GenAI threats and attack techniques, when using LLMs with high security steerability. From a practitioner perspective, ’patching’ a vulnerable application by modifying the system prompt when applicable is almost seamless, eliminating the need to redesign the application and potentially saving hundreds of development hours.

### v4

@@ IV-A Security Steerability Analysis
In Table III, the security steerability of the LLMs is presented alongside their parameter counts. The results reveal a range of policy enforcement scores, varying from 0.379 to 0.837 (Higher is better). Notably, there appears to be a low correlation between the number of parameters and the level of security steerability. This indicates that stronger policy enforcement is not inherently tied to an increase in model size. Instead, achieving better performance in this task requires targeted training to enhance the LLMs’ ability to enforce policies effectively. An evidence for this can be seen in the leading LLM, phi4-mini which has ”only” 3.8B parameters.
TABLE III: Security Steerability Performance Ranking
Model Parameters Security Steerability
phi4-mini:3.8b $3.8\times 10^{9}$ 0.837
falcon3:10b $10.0\times 10^{9}$ 0.833
cogito:8b $8.0\times 10^{9}$ 0.825
llama3.1:8b $8.0\times 10^{9}$ 0.775
gemma2:9b $9.0\times 10^{9}$ 0.767
qwen3:8b $8.0\times 10^{9}$ 0.733
granite3.3:8b $8.0\times 10^{9}$ 0.729
hermes3:8b $8.0\times 10^{9}$ 0.658
granite3.3:2b $2.0\times 10^{9}$ 0.579
falcon3:7b $7.0\times 10^{9}$ 0.567
llama3.2:1b $1.0\times 10^{9}$ 0.521
llama3.2:3b $3.0\times 10^{9}$ 0.504
qwen3:4b $4.0\times 10^{9}$ 0.475
mistral:7b $7.0\times 10^{9}$ 0.458
qwen2.5:7b $7.0\times 10^{9}$ 0.442
qwen2.5:3b $3.0\times 10^{9}$ 0.438
gemma2:2b $2.0\times 10^{9}$ 0.383
hermes3:3b $3.0\times 10^{9}$ 0.379
IV-B Combined Security Assessment
Next, we wanted to perform a comparative analysis on the level of universal security and security steerability. In order to do that, we compared the 240 samples of VeganRibs to sampled version of of 240 attacks from the JailbreakV-28k [3] (24 samples X 10 prohibited content categories). The Figure 1 shows the violation rate (or attack success rate, commonly used in security) of the JailbreakV-28k (blue) compared to VeganRibs (red). The results are connected to lines and ordered from low to high so we can visually see the non-correlation between the two. The analysis demonstrates virtually no correlation between universal security performance and security steerability (Pearson r = 0.125, p = 0.621), confirming that a model’s resistance to ”How to build a bomb” attempts provides no predictive value for its susceptibility to steered harmful outputs through benign-appearing prompts. This pattern is exemplified by extreme cases such as gemma2:9b, which demonstrates poor universal security (80.8% jailbreak violation rate) yet maintains good security steerability (23.3% vegan ribs violation rate, representing a 57.5-percentage-point performance gap), while conversely, smaller models like llama3.2:3b show excellent universal security (2.1% jailbreak violation rate) but poor security steerability (49.6% vegan ribs violation rate).
@@ IV-B Combined Security Assessment
The comparative assessments of VeganRibs and Jailbreak-V28K provide insights into the models’ compliance with their system prompts over user prompts. Models demonstrating high adherence are potential candidates for scenarios where strict policy compliance is critical. Through the evaluation we can witness very different results between security steerability and universal security, although they are both considered ”Security”, and only one of them is traditionally tested. These findings strongly support the conclusion that security steerability represents a distinct, trainable capability that cannot be achieved through conventional scaling or traditional safety training alone, requiring specialized training approaches focused on policy enforcement. Additional breakdown of the results can be seen in Appendix -B.
V Conclusion & Future work
In this work, we highlighted the gap between the conventional notion and benchmarks for LLM universal security and application security - which is the LLM capability to adhere to system instructions in adversarial situations. We named it Security Steerability which represents the first and sole metric known to us for tailoring LLM security to the LLM application, customized according to business logic and specific application threats. We introduced new benchmark dataset comprised with 240 scenarios and attacks dedicated to test security steerability and applied it on 18 leading open-source LLMs.
Our results showed high variability between the LLMs and also challenged the thought that larger LLMs inherently possess better security steerability. Instead, the training methodologies, architectural design, and specialized safety techniques play more significant roles in shaping a model’s ability to enforce policies effectively. This highlights that the relationship between model size and security steerability is non-straightforward, urging a fundamental shift away from scale-focused approaches toward targeted security-oriented training methodologies and architectural innovations.
We then compared the findings with those from commonly utilized LLM security benchmark. We uncovered an intriguing and somewhat unexpected result, that using a benchmark for universal security do not meet the security needs of LLM applications, emphasizing the crucial role of benchmarking and training for security steerability in the creation of reliable and secure GenAI applications.
@@ V Conclusion & Future work
These results should encourage LLM vendors to pay more attention to the security steerability of their models and GenAI application builders to opt for the use of LLMs with higher security steerability. While this study focuses on the LLM security from the perspective of the LLM application security it provides the foundation for research on the effectiveness of system prompt level defenses against GenAI threats and attack techniques, when using LLMs with high security steerability. From a practitioner perspective, ’patching’ a vulnerable application by modifying the system prompt when applicable is almost seamless, eliminating the need to redesign the application and potentially saving hundreds of development hours.